# Non-uniqueness of positive solutions for supercritical semilinear heat equations without scale invariance

### v1

@@ Abstract.
where $N>2$. We assume that the growth rate of $f$ is less than the Joseph-Lundgren exponent for $N>10$ and it satisfies certain assumptions guaranteeing a positive radial singular stationary solution $u^{*}$. We prove that if $u_{0}=u^{*}$, then the problem has at least two positive solutions, namely $u^{*}$ and $u(t)$ which satisfies $u(t)\in L_{loc}^{\infty}(0,t_{0};L^{\infty}(\mathbb{R}^{N}))$ for some $t_{0}>0$ and
@@ 1. Introduction and main results
We define $f(u)=0$ for $u<0$. It is proved in [8, Remark 1.1] that $q_{f}\geq 1$ if the limit $q_{f}$ exists. The growth rate of $f$ can be defined by
@@
and $F^{-1}$ is the inverse function of $F$ defined by
(1.5) $F(u)=\int_{u}^{\infty}\frac{ds}{f(s)}\quad\mbox{for}\quad u>0.$
@@
$L^{p,\infty}(\mathbb{R}^{N}):=\left\{u\in M(\mathbb{R}^{N});\ \left\|u\right\|_{L^{p,\infty}(\mathbb{R}^{N})}:=\sup_{\lambda>0}\left|\left\{x\in\mathbb{R}^{N};\ |u(x)|>\lambda\right\}\right|^{\frac{1}{p}}<\infty\right\},$
@@ 4. Transformations of self-similar solutions
$\int_{\mathbb{R}^{N}}G_{t-s}(x-y)\Delta v(y)\mathop{}\!\mathrm{d}y$ $=\int_{\partial B_{s}}G_{t-s}(x-y)\frac{\partial\bar{u}}{\partial{\nu_{y}}}\mathop{}\!\mathrm{d}S_{y}+\int_{\partial B_{s}^{c}}G_{t-s}(x-y)\frac{\partial u^{*}}{\partial{\nu_{y}}}\mathop{}\!\mathrm{d}S_{y}$
@@
(4.21) $\int_{\delta}^{t-\delta}\int_{\mathbb{R}^{N}}G_{t-s}(x-y)\Delta v\,\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s\geq-\int_{\delta}^{t-\delta}\int_{\mathbb{R}^{N}}\partial_{s}G_{t-s}(x-y)v\,\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s,$
@@
$=\int_{\partial\mathbb{B}_{\delta}}G_{t-s}(x-y)\bar{u}\nu_{1}^{(s)}\mathop{}\!\mathrm{d}S_{y}-\int_{\mathbb{B}_{\delta}}\partial_{s}G_{t-s}(x-y)\bar{u}\mathop{}\!\mathrm{d}(y,s)$
(4.24) $=\int_{\partial\mathbb{B}_{\delta}}G_{t-s}(x-y)\bar{u}\nu_{1}^{(s)}\mathop{}\!\mathrm{d}S_{y}-\int_{\delta}^{t-\delta}\int_{B_{s}}\partial_{s}G_{t-s}(x-y)\bar{u}\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s,$
@@
$\int_{\delta}^{t-\delta}\int_{\mathbb{R}^{N}}$ $G_{t-s}(x-y)v_{s}\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s$
@@
(4.27) $\int_{\delta}^{t-\delta}\int_{\mathbb{R}^{N}}G_{t-s}(x-y)v_{s}\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s\\ =\int_{\delta}^{t-\delta}\int_{\mathbb{R}^{N}}G_{t-s}(x-y)\Delta v\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s+\int_{\delta}^{t-\delta}G_{t-s}(x-y)f(v)\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s.$
@@
$\lim_{\delta\to 0}\int_{\delta}^{t-\delta}\int_{\mathbb{R}^{N}}G_{t-s}(x-y)f(v)\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s=\int_{0}^{t}\int_{\mathbb{R}^{N}}G_{t-s}(x-y)f(v)\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s$
@@
where $\gamma^{*}$ is as in Proposition 1.2 and we used $\lim_{\delta\to 0}\int_{B(0,2)}|v(t-\delta)-v(t)|^{\gamma}\,\mathop{}\!\mathrm{d}x=0$ and Proposition 3.4 with $v(t)\in\mathcal{L}^{{\gamma}}_{ul}(\mathbb{R}^{N})$. Similarly, by Proposition 4.3 we have
@@ 5. Nonuniqueness (supercritical)
$\begin{split}&\nabla S(t)u_{0n}(x)\\ &=\nabla_{x}\left(\int_{\{u^{*}(y)\geq n\}}G_{t}(x-y)n\,\mathop{}\!\mathrm{d}y+\int_{\{u^{*}(y)<n\}}G_{t}(x-y)u^{*}(y)\,\mathop{}\!\mathrm{d}y\right)\\ &=-\int_{\partial\{u^{*}(y)\geq n\}}G_{t}(x-y)n\nu_{1}(y)\,\mathop{}\!\mathrm{d}S(y)\\ &\qquad+\int_{\{u^{*}(y)<n\}}G_{t}(x-y)\nabla u^{*}(y)\,\mathop{}\!\mathrm{d}y-\int_{\partial\{u^{*}(y)<n\}}G_{t}(x-y)u^{*}(y)\nu_{2}(y)\,\mathop{}\!\mathrm{d}S(y)\\ &=\int_{\{u^{*}(y)<n\}}G_{t}(x-y)\nabla u^{*}(y)\,\mathop{}\!\mathrm{d}y\end{split}$
@@
for all $t\in(0,t_{0})$, which implies that $\nabla u^{*}\in L^{\infty}(\mathbb{R}^{N}\times(0,t_{0}))$.

### v2

@@ Abstract.
where $N>2$. We assume that the growth rate of $f$ is less than the Joseph-Lundgren exponent for $N>10$ and it satisfies certain assumptions guaranteeing the existence of a positive radial singular stationary solution $u^{*}$. We prove that if $u_{0}=u^{*}$, then the problem has at least two positive solutions, namely $u^{*}$ and $u(t)$ which satisfies $u(t)\in L_{loc}^{\infty}(0,t_{0};L^{\infty}(\mathbb{R}^{N}))$ for some $t_{0}>0$ and
@@ 1. Introduction and main results
We define $f(u)=0$ for $u<0$. In [8] $A$ is defined by $A=\lim_{u\to\infty}f^{\prime}(u)F(u)$, where
(1.2) $F(u)=\int_{u}^{\infty}\frac{ds}{f(s)}\quad\mbox{for}\quad u>0.$
Then, when (A3) holds, by L’Hospital rule we have
$A=\lim_{u\to\infty}\frac{F(u)}{1/f^{\prime}(u)}=\lim_{u\to\infty}\frac{f^{\prime}(u)^{2}}{f(u)f^{\prime\prime}(u)}=q_{f}.$
It is proved in [8, Remark 1.1] that if the limit $A$ exists, then $A\geq 1$. Hence, $q_{f}\geq 1$ also holds if (A3) holds. The growth rate of $f$ can be defined by
@@
The exponent $q_{S}$ can be formally defined for $N\geq 1$. However, $q_{S}$ becomes effective only in the case $N>2$, since $q_{S}>1$ for $N>2$ and $q_{f}\geq 1$ always holds.
@@
The exponent $q_{JL}$ can be formally defined for $N\geq 1$. However, $q_{JL}$ becomes effective only in the case $N>10$, since $q_{JL}>1$ for $N>10$ and $q_{f}\geq 1$ always holds.
@@
and $F^{-1}$ is the inverse function of $F$ defined by (1.2).
@@
$L^{p,\infty}(\mathbb{R}^{N}):=\left\{u\in M(\mathbb{R}^{N});\ \left\|u\right\|_{L^{p,\infty}(\mathbb{R}^{N})}:=\sup_{\lambda>0}\lambda\left|\left\{x\in\mathbb{R}^{N};\ |u(x)|>\lambda\right\}\right|^{\frac{1}{p}}<\infty\right\},$
@@ 4. Transformations of self-similar solutions
${\int_{B_{s}}G_{t-s}(x-y)\Delta v(y)\mathop{}\!\mathrm{d}y}$ ${+\int_{B^{c}_{s}}G_{t-s}(x-y)\Delta v(y)\mathop{}\!\mathrm{d}y}$
$=\int_{\partial B_{s}}G_{t-s}(x-y)\frac{\partial\bar{u}}{\partial{\nu_{y}}}\mathop{}\!\mathrm{d}S_{y}+\int_{\partial B_{s}^{c}}G_{t-s}(x-y)\frac{\partial u^{*}}{\partial{\nu_{y}}}\mathop{}\!\mathrm{d}S_{y}$
@@
(4.21) ${\int_{\delta}^{t-\delta}\int_{B_{s}}G_{t-s}(x-y)\Delta v\,\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s+\int_{\delta}^{t-\delta}\int_{B_{s}^{c}}G_{t-s}(x-y)\Delta v\,\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s}\\ \geq-\int_{\delta}^{t-\delta}\int_{\mathbb{R}^{N}}\partial_{s}G_{t-s}(x-y)v\,\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s,$
@@
$=\int_{\partial\mathbb{B}_{\delta}}G_{t-s}(x-y)\bar{u}\nu_{1}^{(s)}\mathop{}\!\mathrm{d}S_{{(y,s)}}-\int_{\mathbb{B}_{\delta}}\partial_{s}G_{t-s}(x-y)\bar{u}\mathop{}\!\mathrm{d}(y,s)$
(4.24) $=\int_{\partial\mathbb{B}_{\delta}}G_{t-s}(x-y)\bar{u}\nu_{1}^{(s)}\mathop{}\!\mathrm{d}S_{{(y,s)}}-\int_{\delta}^{t-\delta}\int_{B_{s}}\partial_{s}G_{t-s}(x-y)\bar{u}\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s,$
@@
${\int_{\delta}^{t-\delta}\int_{B_{s}}G_{t-s}(x-y)v_{s}}$ ${\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s+\int_{\delta}^{t-\delta}\int_{B_{s}^{c}}G_{t-s}(x-y)v_{s}\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s}$
@@
(4.27) ${\int_{\delta}^{t-\delta}\int_{B_{s}}G_{t-s}(x-y)v_{s}\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s+\int_{\delta}^{t-\delta}\int_{B^{c}_{s}}G_{t-s}(x-y)v_{s}\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s}\\ ={\int_{\delta}^{t-\delta}\int_{B_{s}}G_{t-s}(x-y)\Delta v\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s+\int_{\delta}^{t-\delta}\int_{B^{c}_{s}}G_{t-s}(x-y)\Delta v\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s}\\ +\int_{\delta}^{t-\delta}{\int_{\mathbb{R}^{N}}}G_{t-s}(x-y)f(v)\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s.$
@@
$\lim_{\delta\to{0^{+}}}\int_{\delta}^{t-\delta}\int_{\mathbb{R}^{N}}G_{t-s}(x-y)f(v)\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s=\int_{0}^{t}\int_{\mathbb{R}^{N}}G_{t-s}(x-y)f(v)\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}s$
@@
where $\gamma^{*}$ is as in Proposition 1.2 and we used $\lim_{\delta\to{0^{+}}}\int_{B(0,2)}|v(t-\delta)-v(t)|^{\gamma}\,\mathop{}\!\mathrm{d}x=0$ and Proposition 3.4 with $v(t)\in\mathcal{L}^{{\gamma}}_{ul}(\mathbb{R}^{N})$. Similarly, by Proposition 4.3 we have
@@ 5. Nonuniqueness (supercritical)
$\begin{split}&\nabla S(t)u_{0n}(x)\\ &=\nabla_{x}\left(\int_{\{u^{*}(y)\geq n\}}G_{t}(x-y)n\,\mathop{}\!\mathrm{d}y+\int_{\{u^{*}(y)<n\}}G_{t}(x-y)u^{*}(y)\,\mathop{}\!\mathrm{d}y\right)\\ &=-\int_{\partial\{u^{*}(y)\geq n\}}G_{t}(x-y)n\nu_{1}(y)\,\mathop{}\!\mathrm{d}{S_{y}}\\ &\qquad+\int_{\{u^{*}(y)<n\}}G_{t}(x-y)\nabla u^{*}(y)\,\mathop{}\!\mathrm{d}y-\int_{\partial\{u^{*}(y)<n\}}G_{t}(x-y)u^{*}(y)\nu_{2}(y)\,\mathop{}\!\mathrm{d}{S_{y}}\\ &=\int_{\{u^{*}(y)<n\}}G_{t}(x-y)\nabla u^{*}(y)\,\mathop{}\!\mathrm{d}y\end{split}$
@@
for all $t\in(0,t_{0})$, which implies that $\nabla u_{n}\in L^{\infty}(\mathbb{R}^{N}\times(0,t_{0}))$.